### v1

# LoRA-PAR: A Flexible Dual-System LoRA Partitioning Approach to Efficient LLM Fine-Tuning

Abstract
Large-scale generative models like DeepSeek-R1 and OpenAI-O1 benefit substantially from chain-of-thought (CoT) reasoning, yet pushing their performance typically requires vast data, large model sizes, and full-parameter fine-tuning. While parameter-efficient fine-tuning (PEFT) helps reduce cost, most existing approaches primarily address domain adaptation or layer-wise allocation rather than explicitly tailoring data and parameters to different response demands. Inspired by “Thinking, Fast and Slow,” which characterizes two distinct modes of thought—System 1 (fast, intuitive, often automatic) and System 2 (slower, more deliberative and analytic)—we draw an analogy that different “subregions” of an LLM’s parameters might similarly specialize for tasks that demand quick, intuitive responses versus those requiring multi-step logical reasoning. Therefore, we propose LoRA-PAR, a dual-system LoRA framework that partitions both data and parameters by System 1 or System 2 demands, using fewer yet more focused parameters for each task. Specifically, we classify task data via multi-model role-playing and voting, and partition parameters based on importance scoring, then adopt a two-stage fine-tuning strategy of training System 1 tasks with supervised fine-tuning (SFT) to enhance knowledge and intuition and refine System 2 tasks with reinforcement learning (RL) to reinforce deeper logical deliberation next. Extensive experiments show that the two-stage fine-tuning strategy, SFT and RL, lowers active parameter usage while matching or surpassing SOTA PEFT baselines.
LoRA-PAR: A Flexible Dual-System LoRA Partitioning Approach to Efficient LLM Fine-Tuning
Yining Huang1††thanks: huangyining1987@gmail.com, Bin Li2††thanks: b.li2@siat.ac.cn, Keke Tang3,4††thanks: tkk2012@gmail.com, Meilian Chen††thanks: 523062863@qq.com 1School of Politics and Public Administration, South China Normal University 2Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 3University of Chinese Academy of Sciences 4Shenyang institute of computing technology, Chinese academy of sciences Correspondence: huangyining1987@gmail.com
1 Introduction
Large language models (LLMs) such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and OpenAI-O1 (Contributors et al., 2024) have shown remarkable progress in complex reasoning when equipped with CoT prompts. However, pushing their performance to new levels often relies on massive datasets and full-parameter fine-tuning, requiring considerable compute and large model sizes. To alleviate this burden, parameter-efficient fine-tuning (PEFT) approaches have emerged as a promising alternative. Most existing PEFT methods, however, predominantly insert uniform adapter modules (e.g., LoRA (Hu et al., 2021)) and do not specifically tailor their parameter configurations to the unique demands of different tasks or reasoning levels. Although there have been some recent attempts to design more task- or data-aware PEFT solutions (Zhang et al., 2023a, b), these works largely focus on domain adaptation or layer-wise parameter allocation rather than explicitly targeting more advanced, multi-step reasoning capabilities.
Meanwhile, inspired by Thinking, Fast and Slow (Kahneman, 2011), we incorporate the dual-system concept into parameter-efficient fine-tuning for LLMs. Specifically, as shown in Figure 2, we draw on the idea that the human brain engages partially distinct neural processes for System 1 versus System 2. Recent studies further evidence that large language models can manifest or benefit from distinct “fast” vs. “slow” modes: (Hagendorff et al., 2022) exhibit human-like intuitive biases, (Pan et al., 2024) propose dynamic decision mechanisms inspired by Kahneman’s framework, and more general discussions support bridging cognitive dual-process theories and AI (Booch et al., 2020). By analogy, we posit that an LLM’s parameters can be partitioned into “subregions” specialized for different response demands. We implement this in three steps: (1) we use multi-expert role-play and voting to classify each training instance into System 1 or System 2 tasks, ensuring quick, direct “fast thinking” problems are kept separate from more deliberative multi-step tasks; (2) we then allocate different subsets of parameters of LoRA modules (via importance-based partitioning) for System 1 and System 2, akin to activating distinct cognitive modes; (3) we train System 1 parameters with end-to-end SFT for direct question–answer mapping, and refine System 2 parameters using reinforcement learning (GRPO (Shao et al., 2024)), similar to how models like DeepSeek-R1 (DeepSeek-AI et al., 2025) achieve deeper chain-of-thought style reasoning. In this way, our approach remains within the lightweight scope of PEFT, while still capturing the dual-process benefits of human cognition—fast, intuitive responses and methodical, step-by-step logic.
2 Related Work
2.1 Parameter Importance Calculation and Pruning
SparseGPT (Frantar and Alistarh, 2023) effectively prunes large-scale LLM parameters without retraining, drastically reducing model size with minimal performance loss. Wanda (Sun et al., 2023) employs activation-aware magnitude pruning without retraining, significantly outperforming traditional magnitude-based methods. LLM-Pruner (Ma et al., 2023) identifies and removes structurally redundant components via gradient-based scoring, retaining general multitask capabilities. Týr-the-Pruner (Li et al., 2025) applies second-order Taylor approximations for global structured pruning, achieving high sparsity levels with minimal accuracy loss.
2.2 Selective Freezing and Dual-Stage Training
LIMA (Zhou et al., 2023) shows minimal fine-tuning effectively aligns pretrained models, implying substantial portions of models can remain frozen without loss of knowledge. ILA (Shi et al., 2024) develops an analysis technique to selectively freeze non-critical layers, improving fine-tuning efficiency and performance. Safety Layer Freezing (Li et al., 2024) suggests freezing identified "safety-critical" layers during further fine-tuning to preserve original alignment and safety behaviors.
2.3 LoRA and PEFT Variants
LoRA (Hu et al., 2021) introduces low-rank adaptation, drastically reducing fine-tuning overhead by freezing most parameters while updating small adapter matrices. PiSSA (Meng et al., 2024) initializes LoRA adapters using pretrained singular vectors, accelerating convergence and boosting task accuracy. OLoRA (Büyükakyüz, 2024) enhances LoRA initialization with orthonormal matrices, significantly accelerating fine-tuning convergence. QLoRA (Dettmers et al., 2023) enables highly efficient 4-bit quantized fine-tuning of large models, drastically lowering computational requirements without performance loss. LoRA+ (Hayou et al., 2024) optimizes LoRA fine-tuning via learning-rate scaling adjustments, achieving faster convergence and higher accuracy.
3 Method
3.1 Overall Workflow
The overall workflow of our proposal proceeds as follows. First, multiple teacher LLMs vote to label each query as fast, single-step (System 1) or multi-step reasoning (System 2). Next, we compute parameter importance in LoRA and keep only the most cumulative-importance score parameters for each system, identifying a shared subset that are important to both. Finally, we apply a two-stage fine-tuning strategy, using SFT for System 1 tasks and RL for System 2. Shared parameters can be partially activated in both stages, controlled by and . This design efficiently addresses “fast vs. slow thinking” within a single LLM by freezing irrelevant parameters and focusing updates on the most crucial subregions.
3.2 Multi-Model Role-Play and Voting for Data Classification
Before partitioning model parameters into different mode of thinking, questions are needed to be identified to fall into which category. Rather than relying on a single classifier—which may be error-prone or biased—we design a multi-model role-playing approach. Here, several advanced LLMs (like the “teachers”) each act to be the “target” model (like the “student”) and classify the questions accordingly. Because these teacher models typically have broader pretraining coverage, they can approximate how the student would perceive the question type—either System 1 or System 2. The prompt of role-playing and example questions of System 1 and System 2 are shown in Figure 3.
As shown in the upper panel of Figure 1 (cf. “Sample Splitter”), each teacher independently provides a classification, and we then apply a voting procedure to aggregate these judgments. This ensures that disagreements—arising from the teachers’ differing architectures or training histories—are resolved in a robust manner. The resulting labeled subsets, (System 1) and (System 2), feed into the subsequent modules, where they guide parameter partitioning and two-stage training.
3.3 Parameter Importance Calculation for Subregion Partitioning
After classifying questions, the next step is to determine which LoRA parameters should be “activated” for each category. We adopt LoRA rather than full-parameter fine-tuning to preserve the base model’s global knowledge and enable a modular strategy of activate or freeze for System 1 and System 2 tasks. The partitioning process parallels how different regions of the human brain are activated in response to different cognitive demands (Kahneman, 2011). In large language models, parameter gradients serve as an analogue to neural activations. If the gradient for a certain parameter is large, it implies that parameter is critical in correcting output errors for a particular task. To improve the model’s ability to answer different types of questions (System 1 or System 2), we apply a mask to ignore the prompt and context tokens in the loss computation—i.e., we focus only on output positions. This ensures our importance scores emphasize each parameter’s contribution to generating the correct final answer rather than merely modeling the prompt text.
Computing Importance Scores.
In practice, we attach LoRA modules at positions of Q/K/V/Gate/Up/Down within the target model layers. Let denote an individual LoRA parameter. We measure its importance via a Taylor expansion of the masked cross-entropy loss up to second order:
Here, is the gradient of the masked loss w.r.t.
, and is the diagonal of the Fisher matrix approximated from per-example gradients . Focusing on the output tokens aligns parameter importance with the model’s ability to produce correct answers.
Selecting and Freezing Parameters.
We rank by and choose the top fraction (controlled by ) as the “activated” subregion for each System. During training, activated parameters remain learnable while the rest are frozen, reducing overhead. Some parameters may appear crucial for both System 1 and System 2; these “overlapping” parameters are shared across the two fine-tuning stages. By partitioning parameters in this manner, our framework moves closer to the neural analogy that different “subregions” are engaged for tasks that demand fast vs. slow thinking.
3.4 Two-Stage Fine-tuning Strategy with Importance-Based Parameter Selection
Building on the importance scores computed in §3.3, we now formalize how to (i) determine how many parameters to activate for each system, (ii) handle the overlap between System 1 and System 2 parameters, and (iii) schedule the fine-tuning process in two distinct stages. As shown in Algorithm 1, our approach hinges on three hyperparameters—, and —that control which parameters and how many are updated for System 1 (SFT) and System 2 (RL).
Threshold : Selecting the Most Important Parameters.
From the parameter-importance visualization (see Figure 4), we observe that System 1 and System 2 each rely on a partially disjoint set of LoRA parameters, with a notable overlap. Moreover, each dataset contains many “low-impact” parameters whose importance is near zero for both systems. We introduce a cumulative-importance threshold . Specifically, for each system’s importance ranking, we keep only the most important subset of parameters with cumulative-importance score over , discarding the tail of negligible-importance parameters to reduce overhead and avoid unnecessary updates. For instance, setting means we retain only the parameters with cumulative-importance score over 90%, they are crucial for System 1 and System 2 tasks, respectively.111The is not the ratio of parameters used in the following training procedure. In the situation of puting LoRA on QKVGUD of LLaMa2 7B, if we set , there are approximately 40% of parameters that will treated as important to System 1/2.
Activation Fractions and : Handling Overlap.
Applying individually to System 1 and System 2 yields two top-ranked sets of LoRA parameters, which partially overlap. Concretely, some parameters rank among the top-ranked for both systems; we call these “shared” (see the purple region in Figure 4). We thus introduce two activation fractions, and , to control how many of these shared parameters are updated during the two training stages:
Stage 1 (SFT on System 1 tasks): We activate (a) all parameters in the System 1-only subset and (b) an fraction of the shared parameters. If , we only partially train the shared region in this stage.
Stage 2 (RL on System 2 tasks): We then activate (a) all parameters in the System 2-only subset and (b) a fraction of the shared parameters. The rest remain frozen, enabling us to flexibly allocate more (or fewer) shared parameters to System 2 based on .
By varying and , we fine-tune the balance between “fast, direct” adaptation for System 1 and “multi-step, deliberative” adaptation for System 2, ensuring that parameters useful for both can be partially or fully trained in each stage as needed.
Why Two Distinct Stages (SFT then RL)?
We adopt SFT RL following practices in OpenAI GPT, DeepSeek-R1, and related literature on multi-stage language model training. System 1 tasks—quick, direct Q&A—are naturally suited to end-to-end SFT, which establishes “fast-thinking” capability without delving into complex reasoning. This “knowledge foundation” helps bootstrap the second stage, where RL encourages step-by-step logical reasoning for System 2 tasks (akin to a “slow-thinking” process). In essence, RL refines and extends the capabilities acquired via SFT, rewarding correct multi-step strategies rather than just direct answers.
Putting It All Together.
Algorithm 1 outlines these steps more formally. In Stage 1 (SFT), only the System 1-only subset plus -portion of shared parameters are trained; in Stage 2 (RL), only the System 2-only subset plus -portion of shared parameters are updated. This design ensures each system’s specialized subregion is honed for its respective tasks, while shared parameters can flexibly contribute to both fast and slow thinking modes.
4 Experiment
4.1 Experimental Setup.
We begin by partitioning each dataset via multi-model role-playing and voting §3.2, then compute LoRA parameter importance and keep the top-ranked for each system §3.3. Training proceeds in two stages §3.4: (1) SFT for System 1, and (2) RL for System 2, with shared parameters managed by and . We measure accuracy across GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021) (trained using Dolly15K (Conover et al., 2023) or OpenPlatypus (Lee et al., 2023)), and HumanEval (Chen et al., 2021) (code tasks), comparing our approach to LoRA (Hu et al., 2021), OLoRA (Büyükakyüz, 2024), PiSSA (Meng et al., 2024), and PiSSA+RL, all based on LLaMA2 7B. Key hyperparameters include (fraction of top-ranked parameters), , (activation fractions for overlapping parameters), and 1–2 training epochs for each baseline.
Data Classification Strategy | Performance |
QwQ w/o role play | 25.32 |
QwQ w/ role play | 26.23 |
Deepseek-R1 w/ role play | 26.84 |
Random partition | 25.85 |
Role play + voting (n=3) | 27.07 |
Role play + voting (n=5) | 27.60 |
4.2 Role-Playing and Voting for Data Classification
We first verify the role-playing and voting approach introduced in §3.2 by comparing various data classification strategies on GSM8K. Specifically, we contrast (a) a single model without role-play, (b) a single model prompted to “act as” LLaMA2 7B, (c) random partitioning, and (d) multiple models with role-play plus voting. As shown in Table 1, the multi-model role-play+voting setup achieves the highest performance. Prompting an external LLM to imitate the target model’s decision boundary (role-playing) reduces misclassification compared to its default inference style, while the voting ensemble mitigates individual biases and yields more robust splits. This result aligns with our intuition that combining multiple “teacher” perspectives better approximates how LLaMA2 7B would distinguish System 1 vs. System 2 questions, ultimately enhancing downstream fine-tuning.
4.3 Adaptive Parameter Usage via
We next investigate how varying the cumulative importance cutpoint (from 0% to 100%) affects both the number of LoRA parameters activated and the resulting performance under SFT. In essence, dictates which and how many parameters are updated, as introduced in §3.4. For each setting, we compare three LoRA module configurations—QKV, GUD, and QKVGUD—against a random-selection baseline that picks the same fraction of parameters but without regard to importance.
Table 3 summarizes our results on GSM8K, and Figure 5 illustrates the overall trend. As increases, performance generally improves, but diminishing returns emerge near the high end. Notably, QKVGUD reaches close to its maximum accuracy even around between 0.8 and 0.9, activating only 30%–40% of LoRA parameters. In contrast, random selection of an equivalent fraction falls short, underscoring that targeting the most important parameters is critical to strong performance. In practice, can thus be tuned to strike a balance between absolute accuracy and parameter budget.
Perf. (SFT) | Perf. (RL) | ||
0 | 0 | 18.88 | 19.03 |
0 | 0.5 | 18.12 | 25.25 |
0 | 1 | 18.35 | 23.58 |
0.5 | 0 | 23.65 | 25.85 |
0.5 | 0.5 | 24.79 | 29.57 |
0.5 | 1 | 24.94 | 31.01 |
1 | 0 | 27.75 | 27.75 |
1 | 0.5 | 26.99 | 31.92 |
1 | 1 | 27.14 | 34.37 |
QKV | GUD | QKVGUD | |||||||
%Param | Perf. | Rand. | %Param | Perf. | Rand. | %Param | Perf. | Rand. | |
0 (Base LLaMA2 7B) | - | 13.3 | - | - | 13.3 | - | - | 13.3 | - |
1 (Original LoRA) | 100 | 26.46 | - | 100 | 30.25 | - | 100 | 30.63 | - |
0.1 | 0.30 | 0 | - | 0.22 | 12.13 | - | 0.32 | 8.87 | - |
0.2 | 1.03 | 1.44 | - | 1.21 | 6.90 | - | 1.22 | 12.59 | - |
0.3 | 2.20 | 15.77 | 0 | 2.89 | 5.76 | 3.11 | 2.72 | 11.90 | 6.29 |
0.4 | 3.93 | 18.44 | - | 5.30 | 6.22 | - | 4.90 | 13.04 | - |
0.5 | 6.39 | 20.55 | - | 8.61 | 11.14 | - | 7.91 | 13.12 | - |
0.6 | 9.88 | 22.29 | 11.22 | 13.10 | 17.36 | 17.13 | 12.07 | 15.16 | 13.87 |
0.7 | 14.90 | 23.73 | - | 19.30 | 21.15 | - | 17.87 | 20.62 | - |
0.8 | 22.52 | 23.65 | - | 28.27 | 22.14 | - | 26.40 | 24.03 | - |
0.9 | 35.70 | 23.35 | 21.30 | 42.90 | 25.47 | 23.43 | 40.56 | 27.3 | 23.43 |
GSM8K | MMLU(Dolly) | MMLU(Platypus) | HumanEval | ||||||
Method | 1 epoch | 2 epoch | 1 epoch | 2 epoch | 1 epoch | 2 epoch | 1 epoch | 2 epoch | |
LLaMa2 7B | 13.3 | – | 46.48 | – | 46.48 | – | 14.39 | – | |
LoRA | LoRA (2021) | 27.8 | 31.86 | 45.85 | 44.99 | 43.36 | 43.16 | 15.37 | 18.54 |
OLoRA (2024) | 29.57 | 32.83 | 45.08 | 45.06 | 45.16 | 45.26 | 15.85 | 19.02 | |
PiSSA (2024) | 30.78 | 33.59 | 22.95 | 23.27 | 24.14 | 23.89 | 21.95 | 24.39 | |
PiSSA+RL | 30.63 | 37.45 | 22.96 | 23.45 | 23.98 | 23.92 | 21.34 | 25.61 | |
Proposal | *PiSSA() | 28.73 | 38.97 | 23.10 | 24.07 | 24.11 | 25.01 | 22.56 | 26.22 |
*PiSSA() | 30.86 | 41.85 | 23.33 | 24.14 | 24.41 | 25.38 | 23.17 | 27.43 | |
*LoRA() | 27.07 | 34.57 | 46.34 | 47.09 | 44.15 | 45.66 | 16.46 | 19.51 |
4.4 Utilizing Shared Parameters via &
Recall that and (introduced in §3.4) control how many of the shared parameters remain active in the SFT (System 1) and RL (System 2) stages, respectively. We fix for each system, then vary and to measure their impact on training dynamics. Table 2 shows results on GSM8K using QKVGUD LoRA. The “Performance (SFT)” column reflects accuracy after the first stage, while “Performance (RL)” is the final accuracy after the second stage. When both , shared parameters remain fully active in both stages—maximizing the SFT checkpoint and yielding the best final score (34.37). Lower values of or reduce the overlap, limiting early gains in SFT or hampering the RL stage’s capacity for multi-step reasoning. In effect, a strong SFT foundation provides a “warm start” for RL (System 2), letting the model build deeper logic on top of its fast-thinking skills.
4.5 Final Performance and Baseline Comparisons
We conclude by evaluating our approach alongside several LoRA-based baselines (LoRA (Hu et al., 2021), OLoRA (Büyükakyüz, 2024), PiSSA (Meng et al., 2024), PiSSA+RL) on four tasks: GSM8K, MMLU (trained with Dolly15K or Platypus), and HumanEval. Each column in Table 4 corresponds to one of these tasks, trained for either one or two epochs. For instance, GSM8K is trained on its own data, whereas MMLU(Dolly) and MMLU(Platypus) use Dolly15K and OpenPlatypus, respectively. HumanEval relies on code-focused data (CodeAlpaca (Chaudhary, 2023), CodeFeedback). The base model (LLaMA2 7B) is shown for reference, with no fine-tuning. Baselines typically perform two rounds of SFT, except for PiSSA+RL, which does one SFT epoch and then a RL epoch. Our proposals (marked with *) apply -based parameter selection ( or ) along with role-play+voting data splits and fully active overlaps () in the two-stage training. Notably, PiSSA( or ) uses about 40% of the full LoRA parameters yet outperforms vanilla PiSSA, indicating that focusing on high-importance subregions can yield stronger results. Overall, our method achieves the best accuracy on GSM8K (41.85%), surpassing PiSSA by about 12% while using significantly fewer parameters. On MMLU, we also observe gains over standard LoRA and PiSSA, confirming that selectively activating only the most relevant parameters is both efficient and effective. Beyond PiSSA, our QKVGUD configuration activates only about 40% of LoRA parameters (based on or ) for each system, yet still outperforms both full LoRA and random subsets of comparable size. As illustrated by the scatter plots in Figure 4, these “top-ranked” parameters form a highly specialized subregion for fast & intuitive (System 1) vs. multi-step (System 2) tasks. In other words, by focusing updates on those parameters that are important to each reasoning style, we realize the dual-system analogy—distinct parameter subsets excel at quick, intuitive SFT or stepwise RL—while lowering active parameter usage.
5 Limitations
Our experiments confirm that selectively activating LoRA parameters for System 1 and System 2 yields clear benefits in both performance and parameter efficiency. By combining role-play-based data splitting with importance-driven parameter partitioning, we effectively approximate the dual-process paradigm within an LLM. Nonetheless, there are a few limitations: (1) Multi-Model Annotations. Although using multiple teacher LLMs improves labeling quality, it adds computational overhead and presupposes access to diverse, high-capacity models. (2) Granularity of Task Partitioning. Our approach treats tasks at a coarse level (System 1 vs. System 2). More nuanced distinctions (e.g., intermediate steps or partial multi-hop reasoning) may require finer-grained analysis. (3) Applicability to Other Architectures. We have shown results on LLaMA2 7B; generalizing to other model families (e.g., decoder-encoder hybrids) may require adjustments in how LoRA parameters are attached and scored.
6 Conclusion
We presented a dual-system PEFT framework inspired by Thinking, Fast and Slow, wherein System 1 and System 2 “subregions” of LoRA parameters address fast, intuitive tasks vs. slower, multi-step reasoning. Our pipeline (i) classifies queries via multi-model role-playing and voting, (ii) determines each LoRA parameter’s importance relative to System 1 or System 2, and (iii) fine-tunes in two stages—SFT for intuitive response, then RL for deeper logic. Across GSM8K, MMLU, and HumanEval, we find that focusing updates on top-ranked parameters not only cuts active parameter usage (often to 40% or less) but also surpasses baseline PEFT methods that uniformly fine-tune larger parameter sets. By assigning each subregion to a distinct “cognitive” mode, we effectively reconcile fast vs. slow thinking within a single LLM. We believe this “subregion specialization” opens new directions for cognitively guided LLM adaptations, enabling more efficient models that still excel at both intuitive and methodical reasoning.